OpenStudy (austinl):

Determine whether the equation is exact. If it is exact, find the solution. \((3x^2-2xy+2)+(6y^2-x^2+3)y\prime=0\)

OpenStudy (anonymous):

basically what you want to do is combine like terms first then you solve from there

OpenStudy (austinl):

You take differential equations?

OpenStudy (anonymous):

Did you check for exactness yes? \[\Large M_y=N_x \]

OpenStudy (austinl):

I haven't worked on this in a while, how exactly would I do that?

OpenStudy (anonymous):

Consider the differential equation in it's original form: \[\Large M(x,y)+N(x,y)y'=0 \] for this differential equation to be exact you must first compute the partial differentials: \[\Large M_y=N_x \] if they are exact you can continue from there by assuming that there exists a function with your desired values.

OpenStudy (anonymous):

That would be any function of the following form: \[\Large F_x+F_yy'=0\] (multivariable chainrule)

OpenStudy (anonymous):

y'=dy/dx write the equation in the formM dx+Ndy=0 then see if